To check the radius of a railroad curve, the effect of a 2-lb weight is observed to be 20.7 lb on a spring scale suspended from the roof of an experimental car rounding the curve at 40mph. what is the radius of the curve?



1
Expert's answer
2021-04-16T07:47:23-0400

"F = \dfrac{mv²}{r}"


"F = 20.7\ Ib× 9.8ms^{-2}"

"F = 202.86\ Ibms^{-2}"


"202.86 = \dfrac{2×40²}{r}"


"r = \dfrac{2× 40²}{202.86}"


"r = 15.77m"